OpenStudy (anonymous):

Two quick questions. Will give medals and will fan. True or Fasle: log (xy) = log(x) log(y) log(xy) = log(x) + log(y)

OpenStudy (anonymous):

A property of Logarithm is this: \[\log(a) + \log(b) = \log(ab)\] So, if one is true then other cannot be.. :)

OpenStudy (anonymous):

false

OpenStudy (anonymous):

true

OpenStudy (anonymous):

There's two different. Which one is false and which is true?

OpenStudy (anonymous):

first 2 second false